OpenStudy (anonymous):

Write one value of a for which the equations of the following system are inconsistent: 2x+y=a 4x =7-2y

OpenStudy (anonymous):

let us write the two equations in slope-intercept form: \[ y=-2x+a\qquad\ldots(1)\\ 2y=-4x+7\implies y=-2x+{7\over2}\qquad\ldots(2) \] now, two equations are inconsistent when their slopes are equal what are the slopes of these two lines?

OpenStudy (anonymous):

-2x?

OpenStudy (anonymous):

just the "-2" part.

OpenStudy (anonymous):

since the slopes are equal, it does not matter what value you choose for "a"; they'd always be in-consistent

OpenStudy (anonymous):

it can be anything besides 7/2 right?

OpenStudy (anonymous):

it can even be 7/2 no limitation cannot be infinite o'course.

OpenStudy (anonymous):

but if it would be 7/2 wouldn't it be consistent?

OpenStudy (anonymous):

how? the definition for in-consistency only deals with the slope. well, if you had a=7/2, then the two lines would be identical. Even then they are inconsistent. But depends on how your teacher taught it. Technically, "a" can be any finite real number.

OpenStudy (anonymous):

oh okay thank you.

OpenStudy (anonymous):

yw